a) Fama's Llamas has a weighted average cost of capital of 12 percent. The company's cost...

a)

Fama's Llamas has a weighted average cost of capital of 12 percent. The company's cost of equity is 12 percent, and its pretax cost of debt is 9 percent. The tax rate is 40 percent. What is the company's target debt-equity ratio?

(Do not round intermediate calculations and round your final answer to 2 decimal places. For example, 1.2345 should be entered as 1.23.)
Answer #1

After-tax cost of debt=9*(1-tax rate)

=9*(1-0.4)=5.4%

Let debt be $x

Equity be $y

Total=$(x+y)

WACC=Respective costs*Respective weight

12=(x/(x+y)*5.4)+(y/(x+y)*12)

12(x+y)=5.4x+12y

12x+12y=5.4x+12y

x=(12-12)y/(12-5.4)

=0

Hence debt-equity ratio=debt/equity

Hence debt-equity ratio=0